OpenStudy (anonymous):

Ok, so I am trying to solve this equation of a line given a point and a slope: m=3/2 (-6,4) I get y=3/2x+13, the answer is supposed to be -3x+2y=26. I am stumped. I used the y-y1=m(x-x1) formula.

OpenStudy (sid1729):

You are correct, your answer is correct but in a different form \[y = \frac{3}{2}x + 13\] Multiply by two on both sides \[2y = 3x + 26\] Subtract 3x from both sides \[-3x + 2y = 26\]

OpenStudy (amistre64):

I hate that formula, never made any sense to me to use a different formula when you got a perfectly good one with y=mx+b You know you "m" and your given your x=-6 and y=4; now solve for "b". 4 = (3/2)(-6) + b 4 = -9 + b 13 = b now throw away your values for y and x and you get y = (3/2)x + 13 -(3/2)x + y = 13 Or if you want integers multiply everything by 2 -3x + 2y = 26

OpenStudy (anonymous):

Thanks sid1729 and amistre64. I think I understand better now. I didn't consider or recognize that the answer I had was the same as the given answer, only in a different form. I don't see why they mess with the y=mx+b form either. It gives ya good practice at manipulating equations, but that is about it, as I see it.